Process Design & Control

Consider the process of filling a tank with compressed air from a constant pressure supply.
Answer #1

Temperature of the tank = 75 oF (534.65 oR)

Volume = 500 ft3

Initial Pressure= 14.7 psia

Using Ideal Gas Law to relate Pressure, Volume and Temperature inside the Tank

\large P*V=n*R*T

n= number moles of air present in the tank

n can be represented as mass of air (m) divided by it molecular weight (mw).

Molecular weight of air is 28.97

mw

.. m = mw * P * R *T

here mw, V, R and T are constant

dm mwV dP

Writing material balance for tank

In = Out + Accumulated

dm 80 * (AP)0.5-0 + dt

The supply of compressed air is at constant pressure of 450 psia

ДР = 450-P

dm -0+ dt 0.5 80 * (450-P)

Substituting dm/dt in above equation

\large \therefore \frac{mw*V}{R*T}*\frac{dP}{dt} = 80*\left ( 450-P \right )^{0.5}

80 R T dt dP V (450 - P)0.5m

Integrating the above equation within limits as

  1. t=0 P =14.7 psia
  2. t=t P=P

R T t dt dP 14.7 (450 - P).5mw* V

80*R* T [t 0] |-2 * (450-P)0.5. - 14.7mw V

80 *R T 0.5 2 * (450-P) + 41.7 = * t

In the above equation, we substitute P=425 psia and calculate time

Time = 1 min